OpenStudy (anonymous):

Find an equation of the line having given slope and containing the given point. m=5/6,(8,-3)

OpenStudy (anonymous):

So this is what you do: -3=5/6(8)+b Solve for b

OpenStudy (anonymous):

-3=20/3+b -20/3 -20/3 -29/3=b

OpenStudy (anonymous):

y=5/6x-29/3